### v1

# Gaussian curvature and Lyapunov exponent as probes of black hole phase transitions

Abstract
We study the Gaussian curvature of unstable null orbits. The Gaussian curvature exhibits multivaluedness near the phase transition point of a first-order phase transition. Numerical investigations of Reissner-Nordström Anti-de Sitter (RN-AdS), Hayward-AdS, and Hayward-Letelier-AdS black holes demonstrate that this geometric multivalued region coincides precisely with the spinodal region calculated by black hole thermodynamics. Using the known relation $K=-\lambda^{2}$ linking orbital geometry to chaotic dynamics, we show that this geometric feature also satisfies the critical exponents predicted by mean-field theory, consistent with those derived from Lyapunov exponents. Our work demonstrates that Gaussian curvature can serve as an alternative effective tool to study the phase structure of black holes.
pacs:
00.00.00
I Introduction
Since the 1970s studies by Bekenstein and Hawking Hawking (1975); Bekenstein (1973), black holes have been regarded as thermodynamic systems. Subsequent studies revealed that black holes can undergo phase transitions, such as the Hawking-Page phase transition Hawking and Page (1983). Later studies showed that when the cosmological constant is treated as pressure, black holes exhibit a van der Waals-like first-order phase transition Kubiznak and Mann (2012) and a reentrant phase transition Altamirano et al. (2013). Furthermore, building on the holographic principle in black hole thermodynamics, Maldacena established the AdS/CFT correspondence Maldacena (1998) in 1997. Further investigation within this duality showed that black holes as quantum thermodynamic systems exhibit chaos and that their Lyapunov exponents obey the MSS bound $\lambda\leq 2\pi T/\hbar$ Shenker and Stanford (2014a, b); Maldacena et al. (2016). In recent years, some studies have applied classical dynamics methods to verify the validity of the MSS bound Hashimoto and Tanahashi (2017); Zhao et al. (2018); Dalui et al. (2019); Lei et al. (2021); Kan and Gwak (2022). The Lyapunov exponent not only directly characterizes the chaotic behavior of black holes as quantum thermodynamic systems, but has also been connected to the imaginary part of quasinormal modes (QNMs) Cardoso et al. (2009); Guo et al. (2022a).
In black hole thermodynamics, the multivaluedness of QNMs Liu et al. (2014); Mahapatra (2016); Chabab et al. (2016); Zou et al. (2017); Zhang et al. (2021); Zhao et al. (2023, 2024, 2025a, 2025b); Hou et al. (2025) and Lyapunov exponents Guo et al. (2022b); Yang et al. (2023); Lyu et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); Karthik R. and Hegde (2025); Awal and Phukon (2025); Yang et al. (2025a) during first-order phase transitions suggest a potential connection between these quantities and black hole phase transitions. Studies of the observable photon sphere have also tied it to black hole phase transitions Wei and Liu (2018); Wei et al. (2019); Zhang et al. (2019); Xu et al. (2019); Li et al. (2020a); Naveena Kumara et al. (2020); Du et al. (2023); Kumar et al. (2024a); Yang et al. (2025b), confirming that dynamical analysis can probe such phenomena. However, one question remains largely unexplored. Since general relativity is fundamentally geometric, how does spacetime geometry itself change during a thermal phase transition? Is there an intrinsic geometric quantity that can directly reflect this transition?
Black holes are exact solutions of Einstein’s equations whose physical processes are reflected by changes in geometric properties. Gaussian curvature, an intrinsic measure of a two-dimensional manifold, directly quantifies local spacetime curvature and serves as an ideal probe of geometry. Using Gaussian curvature, researchers have developed purely geometric methods to study black hole photon spheres Qiao and Li (2022); Chern (2006), thus establishing a clear correspondence between dynamics and geometry. Recently, it has been used to analyze the stability of circular orbits for massless particles, to determine the existence and distribution of stable and unstable orbits Qiao (2025), and to investigate gravitational lensing by massive particles Li et al. (2020b); Qiao and Zhou (2023). A key finding is that the Lyapunov exponent of null circular orbits near black holes is linked to their Gaussian curvature Gallo and Mädler (2025). This raises a previously overlooked question: can geometric quantities such as Gaussian curvature be used to connect with first-order phase transitions of a black hole?
In this work, we investigate unstable particle orbits near a 3+1-dimensional spherically symmetric black hole. We demonstrate that the Gaussian curvature of unstable orbits of massless particles is multivalued at black hole phase transition point, offering a new perspective for understanding black hole phase transitions. This reveals that during such phase transitions, characterized by the existence of multiple spacetime solutions, the corresponding geometric quantity exhibits multivalued behavior.
This paper is arranged as follows. In Sec. II we briefly review the relations of black hole thermodynamics, the Gaussian curvature of two-dimensional surfaces, and the calculation of Lyapunov exponents. In Sec. III we present the investigation of three distinct black holes and the analysis of the results. In Sec. IV we address the computation of critical exponents. In Sec. V we offer a summary and discussion, and the appendix follows. We set $G=c=k_{B}=\hbar=1$ in this paper.
II Theoretical Framework
In this section we first briefly review the relations of black hole thermodynamics, then introduce how to obtain the Gaussian curvature from a two-dimensional Riemannian metric, and finally recall the derivation of the Lyapunov exponent and its connection to the Gaussian curvature.
II.1 Thermodynamics and phase structure of spherically symmetric black holes
In this subsection we briefly review the relations in black hole thermodynamics by considering a 3+1-dimensional spherically symmetric black hole solution
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right).$ (1)
Note that $f(r_{+})=0$, where $r_{+}$ is the horizon radius. The Hawking temperature can be obtained directly from the metric as
$T=\frac{f^{\prime}(r_{+})}{4\pi}.$ (2)
The Gibbs free energy is
$F=M-TS,$ (3)
where $S=A/4=\pi r_{+}^{2}$ is the entropy, $A$ is the area of the black hole horizon. If a first-order phase transition occurs, the free energy exhibits a swallowtail structure, indicating the appearance of three black hole solutions (large black hole, intermediate black hole, and small black hole). The critical point is determined by the following condition
$\frac{\partial T}{\partial r_{+}}=\frac{\partial^{2}T}{\partial r_{+}^{2}}=0.$ (4)
For an Anti-de Sitter spacetime, the pressure is given by Kubiznak and Mann (2012):
$P=-\frac{\Lambda}{8\pi}=\frac{3}{8\pi\ell^{2}},$ (5)
where $\Lambda$ is a cosmological constant and $\ell$ is the AdS radius.
II.2 Gaussian curvature of two-dimensional manifolds
Gaussian curvature $K$ is an intrinsic geometric quantity of the optical metric associated with massless particles, and the light ring is closely linked to black hole spacetime. This suggests that $K$ on the light ring may exhibit anomalous behavior near the phase transition point, offering a potential geometric signature for studying such transitions.
Consider a general 3+1-dimensional spherically symmetric metric
$ds^{2}=-f(r)dt^{2}+\frac{1}{g(r)}dr^{2}+r^{2}d\Omega^{2},$ (6)
where $d\Omega^{2}$ is the unit 2-sphere, and $f(r)$ and $g(r)$ are smooth functions of class $C^{P}$ ($P\geq 2$).
To analyze null geodesics, we rewrite Eq. (6) as the optical metric ($ds^{2}=0$), and restrict our attention to the metric in the equatorial plane ($\theta=\frac{\pi}{2}$)
$dt^{2}=\frac{1}{f(r)}\left(\frac{1}{g(r)}dr^{2}+r^{2}d\phi^{2}\right).$ (7)
Beginning with Eq. (7), which describes a two-dimensional Riemannian manifold, we compute the Gaussian curvature associated with distinct null circular orbits. For a two-dimensional Riemannian manifold in orthogonal coordinates, its Gaussian curvature is given by
$K=-\frac{1}{2\sqrt{EG}}\left\{\frac{\partial}{\partial v}\left[\frac{(E)_{v}}{\sqrt{EG}}\right]+\frac{\partial}{\partial u}\left[\frac{(G)_{u}}{\sqrt{EG}}\right]\right\},$ (8)
where $(u,v)$ are the coordinate variables on the surface, and $E,G$ are the coefficients of the first fundamental form Chern (2006). By setting $E=g_{11}$ and $G=g_{22}$, Eq. (8) translates Gaussian notation into tensor notation
$K=-\frac{1}{2}\frac{1}{\sqrt{g_{rr}g_{\phi\phi}}}\frac{d}{dr}\left(\frac{g^{\prime}_{\phi\phi}}{\sqrt{g_{rr}g_{\phi\phi}}}\right).$ (9)
Substituting Eq. (7) into (9) yields
$\begin{split}K(r)=&-g^{\prime}(r)\frac{2f(r)-rf^{\prime}(r)}{4r}\\ &+\frac{g(r)}{2}\left[f^{\prime}(r)\left(\frac{1}{r}-\frac{f^{\prime}(r)}{f(r)}\right)+f^{\prime\prime}(r)\right].\end{split}$ (10)
For the unstable null circular orbit ($r=r_{LR}$, see the appendix for details), we have the following formula
$f^{\prime}(r_{LR})=\frac{2f(r_{LR})}{r_{LR}}.$ (11)
Inserting Eq. (11) into (10) yields a known relation Gallo and Mädler (2025)
$K(r_{LR})=\left\{\frac{g(r)}{2}\left[f^{\prime\prime}(r)-\frac{f^{\prime}(r)}{r}\right]\right\}\bigg{|}_{r=r_{LR}}.$ (12)
Note that $K(r_{LR})$ depends only on the derivatives of the spherically symmetric black hole metric functions $g(r)$ and $f(r)$. As an intrinsic geometric quantity of a surface, the Gaussian curvature depends solely on the first fundamental form of the surface (a profound result known as the Egregium Theorem). It quantifies the deviation of the first fundamental form of the two-dimensional surface from the Euclidean metric, which is fundamental to the study of intrinsic geometry Chern (2006). As a coordinate-invariant intrinsic curvature, $K$ directly measures spacetime deformation, unlike orbital stability analyses.
In Sec. III we will demonstrate the connection between Gaussian curvature and first-order phase transitions by examining the anomalous behavior of $K$ for circular null orbits near three distinct black holes. A more explicit link between the detailed behavior of $K$ and the phase transition will be provided in the next subsection through the analysis of Lyapunov exponents.
II.3 Lyapunov exponent for unstable orbits
The Lyapunov exponent $\lambda$ is a dynamical quantity that characterizes the chaotic behavior of particle orbits near black holes, and it does not directly reflect a phase transition. In this subsection, we briefly review the derivation of $\lambda$ and its relationship with $K$.
Beginning with the metric Eq. (6), the Lagrangian is
$\mathcal{L}=g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=\sigma,$ (13)
where $\sigma=\{+1,-1,0\}$ denotes timelike, spacelike, and null geodesics, respectively. The angular momentum and energy are defined by
$2L=2\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=2r^{2}\dot{\phi},\quad-2E=2\frac{\partial\mathcal{L}}{\partial\dot{t}}=-2f\dot{t}.$ (14)
On the equatorial plane, the radial motion follows
$(\dot{r})^{2}=V_{\text{eff}}(r),$ (15)
with effective potential
$V_{\text{eff}}(r)=g(r)\left[\frac{E^{2}}{f(r)}-\frac{L^{2}}{r^{2}}-\sigma\right].$ (16)
For massless particles $\sigma=0$, the second derivative on light rings $r_{LR}$ satisfies
$V^{\prime\prime}_{\text{eff}}(r_{LR})=\frac{g(r)L^{2}}{r^{4}f(r)}\left[rf^{\prime}(r)-r^{2}f^{\prime\prime}(r)\right]\bigg{|}_{r=r_{LR}}.$ (17)
Expanding this equation of motion at the light ring yields
$(\dot{\varepsilon})^{2}-\frac{1}{2}V^{\prime\prime}_{\text{eff}}(r_{0})\varepsilon^{2}=0.$ (18)
Note that this is an inverted harmonic oscillator equation, defining the Lyapunov exponent Bhattacharyya et al. (2021); Cardoso et al. (2009)
$\lambda=\sqrt{\frac{1}{2(\dot{t})^{2}}V^{\prime\prime}_{\text{eff}}(r_{0})}.$ (19)
The general solution is
$\varepsilon=Ae^{\lambda t}+Be^{-\lambda t}.$ (20)
Combining Eqs. (14), (19), and (12) yields a known relation Gallo and Mädler (2025)
$K(r_{LR})=-\lambda^{2}(r_{LR}).$ (21)
Eq. (21) establishes a quantitative relation between Gaussian curvature and Lyapunov exponent on the light ring. Previous studies Guo et al. (2022b); Yang et al. (2023); Lyu et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); Karthik R. and Hegde (2025); Awal and Phukon (2025) have demonstrated that $\lambda$ becomes multivalued near the phase transition point. In Refs.  Guo et al. (2022b); Shukla et al. (2024) the authors interpreted the Lyapunov exponent as an order parameter that decreases with increasing temperature $T$. Clearly, Eq. (21) provides a mapping between $K$ and $\lambda$ on the light ring. Such a mapping paves the way for a connection between geometry and thermodynamics, thereby enabling the investigation of thermodynamics from a geometric perspective. Combined with Eq. (21), if $\lambda$ is multivalued in the spinodal region, $K$ must also be multivalued, offering a geometric way to probe black hole phase transitions. Note that not only do null circular orbits admit a Lyapunov exponent, but circular orbits of massive particles can also be defined. In both cases, these dynamical quantities connect with the phase transition.
For completeness, we also examine the Lyapunov exponent associated with the circular orbits of massive particles. The corresponding unstable circular orbit is located at $r=r_{0}$ (see Appendix) with $\sigma=1$, and its Lyapunov exponent is given by
$\begin{split}\lambda&=\frac{1}{2}\left.\sqrt{[2f(r)-rf^{\prime}(r)]V_{\rm eff}^{\prime\prime}(r)}\right|_{r=r_{0}}\\ &=\frac{1}{\sqrt{2}}\sqrt{-\frac{g(r_{0})}{f(r_{0})}\left[\frac{3f(r_{0})f^{\prime}(r_{0})}{r_{0}}-2f^{\prime}(r_{0})^{2}+f(r_{0})f^{\prime\prime}(r_{0})\right]}.\end{split}$ (22)
For $g(r)=f(r)$, the conserved quantities satisfy
$\frac{E}{L}=\frac{\sqrt{f(r_{0})}}{r_{0}}.$ (23)
In Sec. III, we also examine the relationship between $\lambda$ and $\tilde{T}$, where the tilde denotes dimensionless quantities, for massless and massive particles around a specific black hole.
III Numerical Results and Analysis
[…]
III.1 RN-AdS black hole
We begin with the RN-AdS black hole. It is a static, charged, four-dimensional spherically symmetric solution whose metric function is
$f_{R}(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{r^{2}}{\ell^{2}},$ (24)
where $Q$ denotes the black hole’s charge, and $\ell$ is the AdS radius. The Hawking temperature is
$T_{R}=\frac{1}{4\pi r_{+}}\left(1-\frac{Q^{2}}{r_{+}^{2}}+\frac{3r_{+}^{2}}{\ell^{2}}\right).$ (25)
Its Gibbs free energy is given by
$F_{R}=\frac{1}{4}\left(\frac{3Q^{2}}{r_{+}}+r_{+}-\frac{r_{+}^{3}}{\ell^{2}}\right).$ (26)
Here, we use the relation
$M=\frac{r_{+}}{2}\left(1+\frac{Q^{2}}{r_{+}^{2}}+\frac{r_{+}^{2}}{\ell^{2}}\right).$ (27)
By dimensional analysis, we introduce the following scaling
$\tilde{r}_{+}=\frac{r_{+}}{\ell},\quad\tilde{Q}=\frac{Q}{\ell},\quad\tilde{M}=\frac{M}{\ell},\quad\tilde{F}_{R}=\frac{F_{R}}{\ell},\quad\tilde{T}_{R}=T_{R}\ell.$ (28)
The multivalued $\lambda_{R}(\tilde{T}_{R})$ near the phase transition point of RN-AdS black holes is well-established Kubiznak and Mann (2012); Guo et al. (2022b). Here, we present only the behavior of the Gaussian curvature $K_{R}$ for its light rings at $\tilde{T}_{p1}$.
Figure 1: (a) Free energy $\tilde{F}_{R}$ and (b) Gaussian curvature $K_{R}$ of unstable photon orbits versus temperature $\tilde{T}_{R}$ for RN-AdS black holes, and $\tilde{Q}=\frac{1}{8.66},~\tilde{Q}_{c}=\frac{1}{6},~\tilde{Q}<\tilde{Q}_{c}$. Swallowtail structures in $\tilde{F}_{R}$ and multivalued $K_{R}$ at $\tilde{T}_{p1}$ evidence geometric degeneracy during phase transitions.
As shown in Fig. 1, in the canonical ensemble (fixed charge $\tilde{Q}$) and when $\tilde{Q}<\tilde{Q}_{c}$, $\tilde{F}_{R}(\tilde{T}_{R})$ shows the characteristic swallowtail structure, which means a first-order phase transition appears, whose transition point is located at $\tilde{T}_{p1}$. Consequently, in Fig. 1, within the spinodal region $\tilde{T}_{R}\in(\tilde{T}_{R1},\,\tilde{T}_{R2})$, the curve $K_{R}(\tilde{T}_{R})$ exhibits multivalued behavior. This occurs because three black hole solutions (small black hole, intermediate black hole, and large black hole) exist in this region. Consequently, a fixed temperature $\tilde{T}_{R}$ corresponds to multiple values of $K_{R}$, demonstrating that the Gaussian curvature $K_{R}$ is multivalued in the spinodal region.
III.2 Hayward–AdS black hole
Unlike the RN-AdS black hole, the Hayward-AdS black hole is a regular solution in AdS spacetime that avoids the appearance of a spacetime singularity J. (1968); Hayward (2006). Its thermodynamics and phase transitions have been investigated in Ref. Gogoi et al. (2024). The metric function is given by
$f_{H}(r)=1-\frac{2Mr^{2}}{g^{3}+r^{3}}+\frac{r^{2}}{\ell^{2}},$ (29)
where $g$ denotes the magnetic monopole charge arising from the nonlinear electrodynamics term in the action Kumar et al. (2024b), and $\ell$ is the AdS radius. The Hawking temperature is
$T_{H}=\frac{\ell^{2}(r_{+}^{3}-2g^{3})+3r_{+}^{5}}{4\pi\ell^{2}r_{+}(g^{3}+r_{+}^{3})}.$ (30)
The corresponding Gibbs free energy reads
$F_{H}=\frac{2(g^{3}+r_{+}^{3})^{2}\left(1+\frac{r_{+}^{2}}{\ell^{2}}\right)-r_{+}^{3}\left[(r_{+}^{3}-2g^{3})+\frac{3r_{+}^{5}}{\ell^{2}}\right]}{4r_{+}^{2}(g^{3}+r_{+}^{3})}.$ (31)
Here, we use the relation
$M=\left(1+\frac{r_{+}^{2}}{\ell^{2}}\right)\frac{g^{3}+r_{+}^{3}}{2r_{+}^{2}}.$ (32)
We introduce the following scaling
$\tilde{r}_{+}=\frac{r_{+}}{\ell},\quad\tilde{g}=\frac{g}{\ell},\quad\tilde{M}=\frac{M}{\ell},\quad\tilde{F}_{H}=\frac{F_{H}}{\ell},\quad\tilde{T}_{H}=T_{H}\ell.$ (33)
Figure 2: (a) Free energy $\tilde{F}_{H}$ and (b) Gaussian curvature $K_{H}$ of unstable photon orbits versus temperature $\tilde{T}_{H}$ for Hayward-AdS black holes, and $\tilde{g}=0.061538,~\tilde{g}_{c}=0.142336,~\tilde{g}<\tilde{g}_{c}$. Swallowtail structures in $\tilde{F}_{H}$ and multivalued $K_{H}$ at $\tilde{T}_{p2}$ evidence geometric degeneracy during phase transitions.
Previous work has extensively investigated the thermodynamic properties of the Hayward-AdS black hole, including the multivalued behavior of the Lyapunov exponent $\lambda_{H}(\tilde{T}_{H})$ near $\tilde{T}_{p2}$ Gogoi et al. (2024). Here, we focus on presenting the $K_{H}(\tilde{T}_{H})$ curve for null geodesics and its thermodynamic counterpart, the free energy $\tilde{F}_{H}(\tilde{T}_{H})$. Following the same procedure as in the RN-AdS case, the numerical results are shown in Figs. 2.
Figs. 2 and 2 show that, for the Hayward–AdS black hole, when $\tilde{g}<\tilde{g}_{c}$, the Gaussian curvature $K_{H}$ for null geodesics also exhibits multivalued behavior around the first-order phase transition point, as in the RN–AdS case.
III.3 Hayward-Letelier-AdS black hole
[…]
$f_{HL}(r)=1-\frac{2Mr^{2}}{g^{3}+r^{3}}+\frac{r^{2}}{\ell^{2}}-a,$ (34)
where $g$ is the magnetic monopole charge, $\ell$ is the AdS radius, and the string-cloud parameter $a$ arises from the string-cloud term in action Kumar et al. (2024b). When $a=0$, the metric reduces to the Hayward–AdS black hole. The Hawking temperature is
$T_{HL}=\frac{\left[\ell^{2}(r_{+}^{3}-2g^{3})+3r_{+}^{5}+a\ell^{2}(2g^{3}-r_{+}^{3})\right]}{4\pi\ell^{2}r_{+}(r_{+}^{3}+g^{3})}.$ (35)
The Gibbs free energy is
$\begin{split}F_{HL}=&\frac{2(g^{3}+r_{+}^{3})^{2}\left(1-a+\frac{r_{+}^{2}}{\ell^{2}}\right)}{4r_{+}^{2}(g^{3}+r_{+}^{3})}\\ &-\frac{r_{+}^{3}\left[(1-a)(r_{+}^{3}-2g^{3})+\frac{3r_{+}^{5}}{\ell^{2}}\right]}{4r_{+}^{2}(g^{3}+r_{+}^{3})}.\end{split}$ (36)
Here,
$M=\left(1+\frac{r_{+}^{2}}{\ell^{2}}-a\right)\frac{g^{3}+r_{+}^{3}}{2r_{+}^{2}}.$ (37)
Since $a$ is a dimensionless constant, we use the same scaling Eq. (33). Following the same procedure as before, we present the $\tilde{F}_{HL}(\tilde{T}_{HL})$, $K_{HL}(\tilde{T}_{HL})$, and $\lambda_{HL}(\tilde{T}_{HL})$ curves for the Hayward–Letelier–AdS black hole.
(a) $\tilde{F}_{HL}-\tilde{T}_{HL}$
(b) $K_{HL}-\tilde{T}_{HL}$
(c) $\lambda_{HL}-\tilde{T}_{HL}$ (Null)
(d) $\lambda_{HL}-\tilde{T}_{HL}$ (Timelike)
Figure 3: (a) $\tilde{F}_{HL}$, (b) $K_{HL}$ of unstable null orbits and (c-d) $\lambda_{HL}$ of unstable null/timelike orbits versus temperature $\tilde{T}_{HL}$ for Hayward-Letelier-AdS black holes, and $\tilde{g}=0.061538,~\tilde{g}_{c}=0.09002,~a=0.6,~\tilde{g}<\tilde{g}_{c}$. Swallowtail structures in $\tilde{F}_{HL}$ and multivalued $K_{HL}$ at $\tilde{T}_{p3}$ evidence geometric degeneracy during phase transitions. $\lambda_{HL}$ at $\tilde{T}_{p3}$ also exhibits such multivaluedness.
As illustrated in Fig. 3(a), for $\tilde{g}<\tilde{g}_{c}$ the swallowtail structure indicates a first-order phase transition. Similarly, in Figs. 3(b), 3(c), and 3(d), the $K_{HL}(\tilde{T}_{HL})$ curve for null circular orbits and the $\lambda_{HL}(\tilde{T}_{HL})$ curve for both timelike and null orbits exhibit multivalued behavior. This shows that both the Gaussian curvature $K_{HL}$ and the Lyapunov exponent $\lambda_{HL}$ are intimately related to the first-order phase transition, because Eq. (21) connects chaotic dynamics and geometry. This connection paves the way for studying thermodynamics from a geometric perspective.
Moreover, we find that the Gaussian curvature $K_{HL}$ of the unstable null circular orbits in Figs. 12, and 3(b) is consistently negative, in full agreement with the result of Ref. Qiao and Li (2022). In Ref. Qiao and Li (2022), the authors adopted an alternative approach based on the Hadamard theorem, demonstrating that $K<0$ corresponds to unstable circular orbits, while $K>0$ indicates stable ones. Our results show that the Gaussian curvature on the light ring is strictly negative, supporting their conclusion. Additionally, after the black hole undergoes a first-order phase transition, specifically on the right side of the spinodal region, $|K|$ decreases monotonically with increasing temperature $\tilde{T}$. This behavior suggests that $|K|$ may be related to the order parameter, a finding consistent with other recent studies Refs. Guo et al. (2022b); Shukla et al. (2024); Banerjee and Roychowdhury (2012a, b). These results indicate that the Gaussian curvature becomes multivalued whenever the black hole experiences a phase transition.
Figure 4: (a) Free energy $\tilde{F}_{HL}$ and (b) Gaussian curvature $K_{HL}$ versus temperature $\tilde{T}_{HL}$ for a Hayward-Letelier-AdS black hole without phase transition ($\tilde{g}=0.092308,~\tilde{g}_{c}=0.09002,~a=0.6,~\tilde{g}>\tilde{g}_{c}$). Both quantities exhibit monotonic behavior, confirming the absence of thermodynamic criticality.
In contrast, as illustrated in Fig. 4, for the case $\tilde{g}>\tilde{g}_{c}$ where no first-order phase transition occurs, Fig. 4 shows that $K_{HL}(\tilde{T}_{HL})$ becomes a monotonic function of $\tilde{T}_{HL}$. Thus, the multivalued behavior of $K$ and $\lambda$ serves as a robust geometric characterization of the first-order phase transition phenomenon and its associated spinodal region, revealing a profound link between spacetime geometry and thermodynamics.
IV Critical Exponent from Lyapunov Exponents and Gaussian Curvature
Critical exponent characterize the behavior of thermodynamic systems near critical point. This section begins by briefly reviewing how the Lyapunov exponent $\lambda$ of black holes can be used to derive critical exponents for black hole phase transitions using a standard approach proposed in Refs. Banerjee and Roychowdhury (2012a, b). We subsequently examine whether the Gaussian curvature $K$ exhibits an analogous behavior.
Consider the expansion of $\lambda$ near the critical point $\tilde{r}_{c}$
$\lambda_{+}=\lambda_{c}+\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{2},$ (38)
[…]
$\tilde{T}=\tilde{T}_{c}+\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})^{2}+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{3}.$ (39)
This leads to
$\tilde{r}_{+}-\tilde{r}_{c}\approx(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}.$ (40)
Substituting $\tilde{r}_{+}-\tilde{r}_{c}$ into the $\lambda$ expansion yields
$\lambda_{+}-\lambda_{c}=\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}.$ (41)
This implies the following relation:
$\Delta\lambda\sim(\tilde{T}-\tilde{T}_{c})^{\delta_{\lambda}},$ (42)
with the critical exponent $\delta_{\lambda}=1/2$. This result aligns with the analysis for non-spherically symmetric black holes in Refs. Chen et al. (2025); Yang et al. (2025a), and spherically symmetric black holes in Refs. Guo et al. (2022b); Shukla et al. (2024); Lyu et al. (2024). Following their work, the order parameter is defined as $\Delta\lambda\equiv\lambda_{l}-\lambda_{s}$, where $\lambda_{l}$ and $\lambda_{s}$ denote the Lyapunov exponents of large and small black holes, respectively.
We now investigate whether the Gaussian curvature $K$ exhibits similar behavior. We define $\Delta K$ as
$\Delta K\equiv|K(\tilde{r}_{+})|-|K(\tilde{r}_{c})|.$ (43)
Using the relation Eq. (21), we obtain
$\Delta K=\lambda_{+}^{2}-\lambda_{c}^{2}=(\lambda_{+}+\lambda_{c})(\lambda_{+}-\lambda_{c}).$ (44)
Substituting the expression Eq. (41) for $(\lambda_{+}-\lambda_{c})$ yields
$\Delta K=\left(\lambda_{+}+\lambda_{c}\right)\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}.$ (45)
Since $(\lambda_{+}+\lambda_{c})$ is a sum of positive real numbers, i.e. a nonzero constant, we conclude
$\Delta K\sim(\tilde{T}-\tilde{T}_{c})^{\delta_{K}},$ (46)
with identical critical exponent $\delta_{K}=1/2$.
V Discussion
In this work, we systematically investigate the profound connections between spacetime geometry, thermodynamics, and chaos dynamics during first-order phase transitions of black holes. We begin by reviewing the behavior of first-order phase transitions of black holes as thermodynamic systems, noting the limitations of conventional approaches that rely on thermodynamic potentials. We then introduce the Gaussian curvature $K$ as an intrinsic geometric quantity, and using a known relation $K(r_{LR})=-\lambda^{2}(r_{LR})$ for unstable null orbits, we link intrinsic geometry to dynamical chaos and black hole phase transitions.
Through numerical analysis of three spherically symmetric black holes, the RN-AdS, the Hayward-AdS, and the Hayward-Letelier-AdS models, we find that when system parameters $\tilde{Q}$ and $\tilde{g}$ enter the spinodal region, the free energy exhibits a swallowtail structure, while both the Gaussian curvature $K$ and the Lyapunov exponent $\lambda$ display multivalued behavior. The temperature intervals over which this occurs align precisely with the thermodynamic spinodal region. Near the critical point, both $K$ and $\lambda$ diverge with a critical exponent of $1/2$, following the scaling relation $(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}$, consistent with mean-field theory.
Furthermore, we demonstrate that in the absence of a phase transition ($\tilde{g}>\tilde{g}_{c}$), the Gaussian curvature $K$ varies monotonically and the geometric multivaluedness disappears. These results provide the first evidence of spacetime degeneracy accompanying black hole phase transitions and indicate that Gaussian curvature can serve as a diagnostic tool for phase transitions independent of thermodynamic potentials. This opens a new pathway for studying black hole phase structures from a purely geometric perspective.
Appendix A: Potential approach
For a spherically symmetric metric
$ds^{2}=-f(r)dt^{2}+\frac{1}{g(r)}dr^{2}+r^{2}d\Omega^{2},$ (47)
the Lagrangian reads
$\mathcal{L}=g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=\sigma.$ (48)
When $\theta=\frac{\pi}{2}$, we obtain the equation of motion
$(\dot{r})^{2}=V_{\text{eff}}(r).$ (49)
The effective potential is given by
$V_{\text{eff}}(r)=g(r)\left[\frac{E^{2}}{f(r)}-\frac{L^{2}}{r^{2}}-\sigma\right],$ (50)
where $E$ and $L$ are defined by Eq.(14). For massless particles in circular orbits, $\sigma=0$, and the unstable circular orbit is denoted as $r_{LR}$. We can derive from
$V_{\text{eff}}(r_{LR})=0,\quad V^{\prime}_{\text{eff}}(r_{LR})=0.$ (51)
This leads to
$f^{\prime}(r_{LR})=\frac{2f(r_{LR})}{r_{LR}}.$ (52)
For massive particles in circular orbits, $\sigma=1$, and the unstable circular orbit is denoted as $r_{0}$. The orbit satisfies
$V_{\text{eff}}(r_{0})=0,\quad V^{\prime}_{\text{eff}}(r_{0})=0.$ (53)
This results in
$L^{2}=\left(\frac{f^{\prime}r^{3}}{2f-rf^{\prime}}\right)_{r=r_{0}},\quad E^{2}=\left(\frac{2f^{2}}{2f-rf^{\prime}}\right)_{r=r_{0}}.$ (54)
Appendix B: Geometric approach
The light ring $r_{LR}$ can also be determined independently using a geometric approach.
Consider the optical metric given in Eq. (7). Its geodesic curvature $\kappa_{g}$ is given by
$\kappa_{g}=\sqrt{\frac{g(r)}{f(r)}}\frac{2f(r)-rf^{\prime}(r)}{2r}.$ (55)
On the light ring $r=r_{LR}$, the geodesic curvature vanishes, yielding
$2f(r_{LR})=r_{LR}f^{\prime}(r_{LR}).$ (56)
This result coincide with Eq. (52).
VI Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grants Nos. 12473001, 12575049, and 12533001), the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the 111 Project (Grant No. B16009).

### v2

@@
 Abstract
-We study the Gaussian curvature of unstable null orbits.
-The Gaussian curvature exhibits multivaluedness near the phase transition point of a first-order phase transition.
-Numerical investigations of Reissner-Nordström Anti-de Sitter (RN-AdS), Hayward-AdS, and Hayward-Letelier-AdS black holes demonstrate that this geometric multivalued region coincides precisely with the spinodal region calculated by black hole thermodynamics.
-Using the known relation $K=-\lambda^{2}$ linking orbital geometry to chaotic dynamics, we show that this geometric feature also satisfies the critical exponents predicted by mean-field theory, consistent with those derived from Lyapunov exponents.
-Our work demonstrates that Gaussian curvature can serve as an alternative effective tool to study the phase structure of black holes.
+First-order phase transitions of black holes have been extensively studied within thermodynamic frameworks, yet the corresponding evolution of spacetime geometric properties remains unclear.
+This paper establishes a purely differential geometric framework to probe such phase transitions by analyzing the curvature of unstable null orbits.
+Using the geodesic curvature of the optical metric to locate the light ring, we demonstrate that the corresponding Gaussian curvature $K$ serves as a direct geometric signature of the phase transition.
+During a first-order phase transition, the curve $K$ versus temperature $T$ exhibits a multivalued structure within the spinodal region, precisely mirroring the swallowtail behavior of the free energy.
+Numerical analysis of Hayward-Letelier-AdS black holes confirms the effectiveness of this geometric signature.
+Our work demonstrates that the intrinsic geometric quantities of spacetime encode the information of black hole phase transitions.
+These quantities serve simultaneously as geometric probes and order parameters for black hole phase transitions.
+This result provides a purely geometric foundation for understanding the correspondence between thermodynamics and spacetime curvature in the null case.
 pacs:
@@
 (2023); Lyu et al.
+(2024); Kumara et al.
 (2024); Du et al.
@@
 Since general relativity is fundamentally geometric, how does spacetime geometry itself change during a thermal phase transition?
-Is there an intrinsic geometric quantity that can directly reflect this transition?
-Black holes are exact solutions of Einstein’s equations whose physical processes are reflected by changes in geometric properties.
+Is there an intrinsic geometric quantity that can directly reflect this phase transition?
+Black holes are exact solutions of the Einstein equations whose physical processes are reflected by changes in geometric properties.
 Gaussian curvature, an intrinsic measure of a two-dimensional manifold, directly quantifies local spacetime curvature and serves as an ideal probe of geometry.
@@
 (2020b); Qiao and Zhou (2023).
+This raises a previously overlooked question: can geometric quantities such as Gaussian curvature be used to connect with first-order phase transitions of a black hole?
 A key finding is that the Lyapunov exponent of null circular orbits near black holes is linked to their Gaussian curvature Gallo and Mädler (2025).
-This raises a previously overlooked question: can geometric quantities such as Gaussian curvature be used to connect with first-order phase transitions of a black hole?
+The established role of the Lyapunov exponent as an effective probe makes it possible to address the above question.
 In this work, we investigate unstable particle orbits near a 3+1-dimensional spherically symmetric black hole.
@@
 In Sec.
-II we briefly review the relations of black hole thermodynamics, the Gaussian curvature of two-dimensional surfaces, and the calculation of Lyapunov exponents.
+II, we briefly review the relations of black hole thermodynamics, the Gaussian curvature of two-dimensional surfaces, and the calculation of Lyapunov exponents.
+In Sect.
+III, we present the investigation of the Hayward-Letelier-AdS black hole and the analysis of the results.
 In Sec.
-III we present the investigation of three distinct black holes and the analysis of the results.
+IV, we address the computation of critical exponents.
 In Sec.
-IV we address the computation of critical exponents.
-In Sec.
-V we offer a summary and discussion, and the appendix follows.
+V, we offer a summary and discussion, and the appendix follows.
 We set $G=c=k_{B}=\hbar=1$ in this paper.
 II Theoretical Framework
-In this section we first briefly review the relations of black hole thermodynamics, then introduce how to obtain the Gaussian curvature from a two-dimensional Riemannian metric, and finally recall the derivation of the Lyapunov exponent and its connection to the Gaussian curvature.
+In this section, we first briefly review the relations of black hole thermodynamics, then introduce how to obtain the Gaussian curvature from a two-dimensional Riemannian metric, and finally recall the derivation of the Lyapunov exponent and its connection to the Gaussian curvature.
 II.1 Thermodynamics and phase structure of spherically symmetric black holes
-In this subsection we briefly review the relations in black hole thermodynamics by considering a 3+1-dimensional spherically symmetric black hole solution
+In this subsection, we briefly review the relations in black hole thermodynamics by considering a 3+1-dimensional spherically symmetric black hole solution
 $ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right).$ (1)
-Note that $f(r_{+})=0$, where $r_{+}$ is the horizon radius. The Hawking temperature can be obtained directly from the metric as
-$T=\frac{f^{\prime}(r_{+})}{4\pi}.$ (2)
-The Gibbs free energy is
+Note that $f(r_{+})=0$, where $r_{+}$ is the horizon radius.
+The temperature can be obtained directly from the first law of a charged and static black hole in extended phase space as Kumar et al.
+(2024b)
+$T=\frac{\partial M}{\partial S},$ (2)
+for a fixed charge and pressure. The Gibbs free energy is
 $F=M-TS,$ (3)
@@
 Gaussian curvature $K$ is an intrinsic geometric quantity of the optical metric associated with massless particles, and the light ring is closely linked to black hole spacetime.
-This suggests that $K$ on the light ring may exhibit anomalous behavior near the phase transition point, offering a potential geometric signature for studying such transitions.
-Consider a general 3+1-dimensional spherically symmetric metric
+This suggests that $K$ on the light ring may exhibit anomalous behavior near the phase transition point, offering a potential geometric signature for studying such phase transitions.
+Consider a static 3+1-dimensional spherically symmetric metric
 $ds^{2}=-f(r)dt^{2}+\frac{1}{g(r)}dr^{2}+r^{2}d\Omega^{2},$ (6)
@@
 $\begin{split}K(r)=&-g^{\prime}(r)\frac{2f(r)-rf^{\prime}(r)}{4r}\\ &+\frac{g(r)}{2}\left[f^{\prime}(r)\left(\frac{1}{r}-\frac{f^{\prime}(r)}{f(r)}\right)+f^{\prime\prime}(r)\right].\end{split}$ (10)
-For the unstable null circular orbit ($r=r_{LR}$, see the appendix for details), we have the following formula
-$f^{\prime}(r_{LR})=\frac{2f(r_{LR})}{r_{LR}}.$ (11)
-Inserting Eq. (11) into (10) yields a known relation Gallo and Mädler (2025)
+Consider the optical metric given in Eq. (7), its geodesic curvature $\kappa_{g}$ is given by
+$\kappa_{g}=\sqrt{\frac{g(r)}{f(r)}}\frac{2f(r)-rf^{\prime}(r)}{2r}.$ (11)
+On the unstable null circular orbit $r=r_{LR}$ (the light ring), the geodesic curvature vanishes, yielding
+$2f(r_{LR})=r_{LR}f^{\prime}(r_{LR}).$ (12)
+For the unstable null circular orbit ($r=r_{LR}$, inserting Eq. (12) into (10) yields a known relation Gallo and Mädler (2025)
 $K(r_{LR})=\left\{\frac{g(r)}{2}\left[f^{\prime\prime}(r)-\frac{f^{\prime}(r)}{r}\right]\right\}\bigg|_{r=r_{LR}}.$ (13)
@@
 In Sec.
-III we will demonstrate the connection between Gaussian curvature and first-order phase transitions by examining the anomalous behavior of $K$ for circular null orbits near three distinct black holes.
+III, we will demonstrate the connection between Gaussian curvature and first-order phase transitions by examining the anomalous behavior of $K$ for circular null orbits near the Hayward-Letelier-AdS.
 A more explicit link between the detailed behavior of $K$ and the phase transition will be provided in the next subsection through the analysis of Lyapunov exponents.
@@
 $\mathcal{L}=g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=\sigma,$ (14)
-where $\sigma=\{+1,-1,0\}$ denotes timelike, spacelike, and null geodesics, respectively. The angular momentum and energy are defined by
+where $\sigma=\{+1,-1,0\}$ denotes spacelike, timelike, and null geodesics, respectively. The angular momentum and energy are defined by
 $2L=2\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=2r^{2}\dot{\phi},\quad-2E=2\frac{\partial\mathcal{L}}{\partial\dot{t}}=-2f\dot{t}.$ (15)
@@
 Expanding this equation of motion at the light ring yields
-$(\dot{\varepsilon})^{2}-\frac{1}{2}V^{\prime\prime}_{\text{eff}}(r_{0})\varepsilon^{2}=0.$ (18)
-Note that this is an inverted harmonic oscillator equation, defining the Lyapunov exponent Bhattacharyya et al. (2021); Cardoso et al. (2009)
-$\lambda=\sqrt{\frac{1}{2(\dot{t})^{2}}V^{\prime\prime}_{\text{eff}}(r_{0})}.$ (19)
+$(\dot{\varepsilon})^{2}-\frac{1}{2}V^{\prime\prime}_{\text{eff}}(r_{LR})\varepsilon^{2}=0,$ (19)
+where $\varepsilon=r-r_{LR}$. Note that this is an inverted harmonic oscillator equation, defining the Lyapunov exponent Bhattacharyya et al. (2021); Cardoso et al. (2009)
+$\lambda=\sqrt{\frac{1}{2(\dot{t})^{2}}V^{\prime\prime}_{\text{eff}}(r_{LR})}.$ (20)
 The general solution is
@@
 (2022b); Shukla et al.
-(2024) the authors interpreted the Lyapunov exponent as an order parameter that decreases with increasing temperature $T$.
+(2024), the authors interpreted the Lyapunov exponent as an order parameter that decreases with increasing temperature $T$.
 Clearly, Eq.
@@
 For completeness, we also examine the Lyapunov exponent associated with the circular orbits of massive particles.
-The corresponding unstable circular orbit is located at $r=r_{0}$ (see Appendix) with $\sigma=1$, and its Lyapunov exponent is given by
+The corresponding unstable circular orbit is located at $r=r_{0}$ with $\sigma=-1$, and its Lyapunov exponent is given by
 $\begin{split}\lambda&=\frac{1}{2}\left.\sqrt{[2f(r)-rf^{\prime}(r)]V_{\rm eff}^{\prime\prime}(r)}\right|_{r=r_{0}}\\ &=\frac{1}{\sqrt{2}}\sqrt{-\frac{g(r_{0})}{f(r_{0})}\left[\frac{3f(r_{0})f^{\prime}(r_{0})}{r_{0}}-2f^{\prime}(r_{0})^{2}+f(r_{0})f^{\prime\prime}(r_{0})\right]}.\end{split}$ (23)
@@
 […]
-III.1 RN-AdS black hole
-We begin with the RN-AdS black hole. It is a static, charged, four-dimensional spherically symmetric solution whose metric function is
-$f_{R}(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{r^{2}}{\ell^{2}},$ (24)
-where $Q$ denotes the black hole’s charge, and $\ell$ is the AdS radius. The Hawking temperature is
-$T_{R}=\frac{1}{4\pi r_{+}}\left(1-\frac{Q^{2}}{r_{+}^{2}}+\frac{3r_{+}^{2}}{\ell^{2}}\right).$ (25)
-Its Gibbs free energy is given by
-$F_{R}=\frac{1}{4}\left(\frac{3Q^{2}}{r_{+}}+r_{+}-\frac{r_{+}^{3}}{\ell^{2}}\right).$ (26)
-Here, we use the relation
-$M=\frac{r_{+}}{2}\left(1+\frac{Q^{2}}{r_{+}^{2}}+\frac{r_{+}^{2}}{\ell^{2}}\right).$ (27)
-By dimensional analysis, we introduce the following scaling
-$\tilde{r}_{+}=\frac{r_{+}}{\ell},\quad\tilde{Q}=\frac{Q}{\ell},\quad\tilde{M}=\frac{M}{\ell},\quad\tilde{F}_{R}=\frac{F_{R}}{\ell},\quad\tilde{T}_{R}=T_{R}\ell.$ (28)
-The multivalued $\lambda_{R}(\tilde{T}_{R})$ near the phase transition point of RN-AdS black holes is well-established Kubiznak and Mann (2012); Guo et al.
-(2022b).
-Here, we present only the behavior of the Gaussian curvature $K_{R}$ for its light rings at $\tilde{T}_{p1}$.
-Figure 1: (a) Free energy $\tilde{F}_{R}$ and (b) Gaussian curvature $K_{R}$ of unstable photon orbits versus temperature $\tilde{T}_{R}$ for RN-AdS black holes, and $\tilde{Q}=\frac{1}{8.66},~\tilde{Q}_{c}=\frac{1}{6},~\tilde{Q}<\tilde{Q}_{c}$.
-Swallowtail structures in $\tilde{F}_{R}$ and multivalued $K_{R}$ at $\tilde{T}_{p1}$ evidence geometric degeneracy during phase transitions.
-As shown in Fig.
-1, in the canonical ensemble (fixed charge $\tilde{Q}$) and when $\tilde{Q}<\tilde{Q}_{c}$, $\tilde{F}_{R}(\tilde{T}_{R})$ shows the characteristic swallowtail structure, which means a first-order phase transition appears, whose transition point is located at $\tilde{T}_{p1}$.
-Consequently, in Fig.
-1, within the spinodal region $\tilde{T}_{R}\in(\tilde{T}_{R1},\,\tilde{T}_{R2})$, the curve $K_{R}(\tilde{T}_{R})$ exhibits multivalued behavior.
-This occurs because three black hole solutions (small black hole, intermediate black hole, and large black hole) exist in this region.
-Consequently, a fixed temperature $\tilde{T}_{R}$ corresponds to multiple values of $K_{R}$, demonstrating that the Gaussian curvature $K_{R}$ is multivalued in the spinodal region.
-III.2 Hayward–AdS black hole
-Unlike the RN-AdS black hole, the Hayward-AdS black hole is a regular solution in AdS spacetime that avoids the appearance of a spacetime singularity J.
-(1968); Hayward (2006).
-Its thermodynamics and phase transitions have been investigated in Ref.
-Gogoi et al.
-(2024).
-The metric function is given by
-$f_{H}(r)=1-\frac{2Mr^{2}}{g^{3}+r^{3}}+\frac{r^{2}}{\ell^{2}},$ (29)
-where $g$ denotes the magnetic monopole charge arising from the nonlinear electrodynamics term in the action Kumar et al. (2024b), and $\ell$ is the AdS radius. The Hawking temperature is
-$T_{H}=\frac{\ell^{2}(r_{+}^{3}-2g^{3})+3r_{+}^{5}}{4\pi\ell^{2}r_{+}(g^{3}+r_{+}^{3})}.$ (30)
-The corresponding Gibbs free energy reads
-$F_{H}=\frac{2(g^{3}+r_{+}^{3})^{2}\left(1+\frac{r_{+}^{2}}{\ell^{2}}\right)-r_{+}^{3}\left[(r_{+}^{3}-2g^{3})+\frac{3r_{+}^{5}}{\ell^{2}}\right]}{4r_{+}^{2}(g^{3}+r_{+}^{3})}.$ (31)
-Here, we use the relation
-$M=\left(1+\frac{r_{+}^{2}}{\ell^{2}}\right)\frac{g^{3}+r_{+}^{3}}{2r_{+}^{2}}.$ (32)
-We introduce the following scaling
-$\tilde{r}_{+}=\frac{r_{+}}{\ell},\quad\tilde{g}=\frac{g}{\ell},\quad\tilde{M}=\frac{M}{\ell},\quad\tilde{F}_{H}=\frac{F_{H}}{\ell},\quad\tilde{T}_{H}=T_{H}\ell.$ (33)
-Figure 2: (a) Free energy $\tilde{F}_{H}$ and (b) Gaussian curvature $K_{H}$ of unstable photon orbits versus temperature $\tilde{T}_{H}$ for Hayward-AdS black holes, and $\tilde{g}=0.061538,~\tilde{g}_{c}=0.142336,~\tilde{g}<\tilde{g}_{c}$.
-Swallowtail structures in $\tilde{F}_{H}$ and multivalued $K_{H}$ at $\tilde{T}_{p2}$ evidence geometric degeneracy during phase transitions.
-Previous work has extensively investigated the thermodynamic properties of the Hayward-AdS black hole, including the multivalued behavior of the Lyapunov exponent $\lambda_{H}(\tilde{T}_{H})$ near $\tilde{T}_{p2}$ Gogoi et al.
-(2024).
-Here, we focus on presenting the $K_{H}(\tilde{T}_{H})$ curve for null geodesics and its thermodynamic counterpart, the free energy $\tilde{F}_{H}(\tilde{T}_{H})$.
-Following the same procedure as in the RN-AdS case, the numerical results are shown in Figs.
-2.
-Figs.
-2 and 2 show that, for the Hayward–AdS black hole, when $\tilde{g}<\tilde{g}_{c}$, the Gaussian curvature $K_{H}$ for null geodesics also exhibits multivalued behavior around the first-order phase transition point, as in the RN–AdS case.
-III.3 Hayward-Letelier-AdS black hole
 […]
@@
 (2024b).
+For studies of Hayward-AdS black holes with string fluids in the extended thermodynamic phase space, see Ref.
+Anand et al.
+(2025) for thermodynamic curvature and topological properties, and Ref.
+Singh et al.
+(2025) for chaotic behavior.
 When $a=0$, the metric reduces to the Hayward–AdS black hole.
-The Hawking temperature is
-$T_{HL}=\frac{\left[\ell^{2}(r_{+}^{3}-2g^{3})+3r_{+}^{5}+a\ell^{2}(2g^{3}-r_{+}^{3})\right]}{4\pi\ell^{2}r_{+}(r_{+}^{3}+g^{3})}.$ (35)
+The temperature is Anand et al.
+(2025)
+$T_{HL}=\frac{r_{+}^{3}\left[(1-a)+3r_{+}^{2}/\ell^{2}\right]-2g^{3}(1-a)}{4\pi r_{+}^{4}}.$ (26)
 The Gibbs free energy is
-$\begin{split}F_{HL}=&\frac{2(g^{3}+r_{+}^{3})^{2}\left(1-a+\frac{r_{+}^{2}}{\ell^{2}}\right)}{4r_{+}^{2}(g^{3}+r_{+}^{3})}\\ &-\frac{r_{+}^{3}\left[(1-a)(r_{+}^{3}-2g^{3})+\frac{3r_{+}^{5}}{\ell^{2}}\right]}{4r_{+}^{2}(g^{3}+r_{+}^{3})}.\end{split}$ (36)
+$\begin{split}F_{HL}=\frac{1}{4}\left\{g^{3}\left[\frac{2}{\ell^{2}}+\frac{4(1-a)}{r_{+}^{2}}\right]+(1-a)r_{+}-\frac{r_{+}^{3}}{\ell^{2}}\right\}.\end{split}$ (27)
 Here,
 $M=\left(1+\frac{r_{+}^{2}}{\ell^{2}}-a\right)\frac{g^{3}+r_{+}^{3}}{2r_{+}^{2}}.$ (28)
-Since $a$ is a dimensionless constant, we use the same scaling Eq.
-(33).
-Following the same procedure as before, we present the $\tilde{F}_{HL}(\tilde{T}_{HL})$, $K_{HL}(\tilde{T}_{HL})$, and $\lambda_{HL}(\tilde{T}_{HL})$ curves for the Hayward–Letelier–AdS black hole.
+Since $a$ is a dimensionless constant, We introduce the following scaling
+$\tilde{r}_{+}=\frac{r_{+}}{\ell},\quad\tilde{g}=\frac{g}{\ell},\quad\tilde{M}=\frac{M}{\ell},\quad\tilde{F}_{H}=\frac{F_{H}}{\ell},\quad\tilde{T}_{H}=T_{H}\ell.$ (29)
+We present the $\tilde{F}_{HL}(\tilde{T}_{HL})$, $K_{HL}(\tilde{T}_{HL})$, and $\lambda_{HL}(\tilde{T}_{HL})$ curves for the Hayward–Letelier–AdS black hole.
 (a) $\tilde{F}_{HL}-\tilde{T}_{HL}$
@@
 (d) $\lambda_{HL}-\tilde{T}_{HL}$ (Timelike)
-Figure 3: (a) $\tilde{F}_{HL}$, (b) $K_{HL}$ of unstable null orbits and (c-d) $\lambda_{HL}$ of unstable null/timelike orbits versus temperature $\tilde{T}_{HL}$ for Hayward-Letelier-AdS black holes, and $\tilde{g}=0.061538,~\tilde{g}_{c}=0.09002,~a=0.6,~\tilde{g}<\tilde{g}_{c}$.
-Swallowtail structures in $\tilde{F}_{HL}$ and multivalued $K_{HL}$ at $\tilde{T}_{p3}$ evidence geometric degeneracy during phase transitions.
-$\lambda_{HL}$ at $\tilde{T}_{p3}$ also exhibits such multivaluedness.
+Figure 1: Thermodynamic, geometric and chaos signatures of the first-order phase transition in Hayward-Letelier-AdS space-times.
+(a) $\tilde{F}_{HL}$, (b) $K_{HL}$ of unstable null orbits and (c-d) $\lambda_{HL}$ of unstable null/timelike orbits versus temperature $\tilde{T}_{HL}$ for Hayward-Letelier-AdS black holes, and $\tilde{g}=0.0615,~\tilde{g}_{c}=0.1042,~a=0.6,~\tilde{g}<\tilde{g}_{c},~L=20\ell$.
+Swallowtail structures in $\tilde{F}_{HL}$ and multivalued $K_{HL}$ at $\tilde{T}_{p}$ evidence geometric degeneracy during phase transitions.
+$\lambda_{HL}$ at $\tilde{T}_{p}$ also exhibits such multivaluedness.
 As illustrated in Fig.
-3(a), for $\tilde{g}<\tilde{g}_{c}$ the swallowtail structure indicates a first-order phase transition.
+1(a), for $\tilde{g}<\tilde{g}_{c}$ the swallowtail structure indicates a first-order phase transition.
 Similarly, in Figs.
-3(b), 3(c), and 3(d), the $K_{HL}(\tilde{T}_{HL})$ curve for null circular orbits and the $\lambda_{HL}(\tilde{T}_{HL})$ curve for both timelike and null orbits exhibit multivalued behavior.
+1(b), 1(c), and 1(d), the $K_{HL}(\tilde{T}_{HL})$ curve for null circular orbits and the $\lambda_{HL}(\tilde{T}_{HL})$ curve for both timelike and null orbits exhibit multivalued behavior.
 This shows that both the Gaussian curvature $K_{HL}$ and the Lyapunov exponent $\lambda_{HL}$ are intimately related to the first-order phase transition, because Eq.
@@
 This connection paves the way for studying thermodynamics from a geometric perspective.
-Moreover, we find that the Gaussian curvature $K_{HL}$ of the unstable null circular orbits in Figs.
-12, and 3(b) is consistently negative, in full agreement with the result of Ref.
+Moreover, we find that the Gaussian curvature $K_{HL}$ of the unstable null circular orbits in Fig.
+1(b) is consistently negative, in full agreement with the result of Ref.
 Qiao and Li (2022).
@@
 Additionally, after the black hole undergoes a first-order phase transition, specifically on the right side of the spinodal region, $|K|$ decreases monotonically with increasing temperature $\tilde{T}$.
-This behavior suggests that $|K|$ may be related to the order parameter, a finding consistent with other recent studies Refs.
+This behavior suggests that $K$ may be related to the order parameter, a finding consistent with other recent studies Refs.
 Guo et al.
@@
 These results indicate that the Gaussian curvature becomes multivalued whenever the black hole experiences a phase transition.
-Figure 4: (a) Free energy $\tilde{F}_{HL}$ and (b) Gaussian curvature $K_{HL}$ versus temperature $\tilde{T}_{HL}$ for a Hayward-Letelier-AdS black hole without phase transition ($\tilde{g}=0.092308,~\tilde{g}_{c}=0.09002,~a=0.6,~\tilde{g}>\tilde{g}_{c}$).
+Figure 2: Thermodynamic and geometric signatures of the absence of a first-order phase transition in Hayward-Letelier-AdS space-times.
+(a) Free energy $\tilde{F}_{HL}$ and (b) Gaussian curvature $K_{HL}$ versus temperature $\tilde{T}_{HL}$ for a Hayward-Letelier-AdS black hole without phase transition ($\tilde{g}=0.0615,~\tilde{g}_{c}=0.1042,~a=0.6,~\tilde{g}>\tilde{g}_{c}$).
 Both quantities exhibit monotonic behavior, confirming the absence of thermodynamic criticality.
 In contrast, as illustrated in Fig.
-4, for the case $\tilde{g}>\tilde{g}_{c}$ where no first-order phase transition occurs, Fig.
+2, for the case $\tilde{g}>\tilde{g}_{c}$ where no first-order phase transition occurs, Fig.
 2 shows that $K_{HL}(\tilde{T}_{HL})$ becomes a monotonic function of $\tilde{T}_{HL}$.
@@
 IV Critical Exponent from Lyapunov Exponents and Gaussian Curvature
-Critical exponent characterize the behavior of thermodynamic systems near critical point.
-This section begins by briefly reviewing how the Lyapunov exponent $\lambda$ of black holes can be used to derive critical exponents for black hole phase transitions using a standard approach proposed in Refs.
-Banerjee and Roychowdhury (2012a, b).
-We subsequently examine whether the Gaussian curvature $K$ exhibits an analogous behavior.
-Consider the expansion of $\lambda$ near the critical point $\tilde{r}_{c}$
-$\lambda_{+}=\lambda_{c}+\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{2},$ (38)
-[…]
-$\tilde{T}=\tilde{T}_{c}+\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})^{2}+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{3}.$ (39)
-This leads to
-$\tilde{r}_{+}-\tilde{r}_{c}\approx(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}.$ (40)
-Substituting $\tilde{r}_{+}-\tilde{r}_{c}$ into the $\lambda$ expansion yields
-$\lambda_{+}-\lambda_{c}=\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}.$ (41)
-This implies the following relation:
-$\Delta\lambda\sim(\tilde{T}-\tilde{T}_{c})^{\delta_{\lambda}},$ (42)
-with the critical exponent $\delta_{\lambda}=1/2$.
-This result aligns with the analysis for non-spherically symmetric black holes in Refs.
-Chen et al.
-(2025); Yang et al.
-(2025a), and spherically symmetric black holes in Refs.
-Guo et al.
-(2022b); Shukla et al.
-(2024); Lyu et al.
-(2024).
-Following their work, the order parameter is defined as $\Delta\lambda\equiv\lambda_{l}-\lambda_{s}$, where $\lambda_{l}$ and $\lambda_{s}$ denote the Lyapunov exponents of large and small black holes, respectively.
-We now investigate whether the Gaussian curvature $K$ exhibits similar behavior. We define $\Delta K$ as
-$\Delta K\equiv|K(\tilde{r}_{+})|-|K(\tilde{r}_{c})|.$ (43)
-Using the relation Eq. (21), we obtain
-$\Delta K=\lambda_{+}^{2}-\lambda_{c}^{2}=(\lambda_{+}+\lambda_{c})(\lambda_{+}-\lambda_{c}).$ (44)
-Substituting the expression Eq. (41) for $(\lambda_{+}-\lambda_{c})$ yields
-$\Delta K=\left(\lambda_{+}+\lambda_{c}\right)\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}.$ (45)
-Since $(\lambda_{+}+\lambda_{c})$ is a sum of positive real numbers, i.e. a nonzero constant, we conclude
-$\Delta K\sim(\tilde{T}-\tilde{T}_{c})^{\delta_{K}},$ (46)
-with identical critical exponent $\delta_{K}=1/2$.
+Critical exponents are used to describe the behavior of physical quantities of thermodynamic systems near the critical point and serve as important physical quantities for characterizing phase transitions.
+At the phase transition point, the small and large black hole branches coexist, leading to discontinuities in both the null Lyapunov exponent and the Gaussian curvature.
+This section will examine the continuity of the null Lyapunov exponent and Gaussian curvature for the Hayward–Letelier–AdS black hole at the phase transition point $T_{p}$, as well as their discontinuity near $T_{p}$, investigating whether they can serve as order parameters and calculating their critical exponents.
+We define the null Lyapunov exponents for the small and large black hole branches within the spinodal region as $\lambda_{S}$ and $\lambda_{L}$, respectively, and the corresponding Gaussian curvatures as $K_{S}$ and $K_{L}$.
+Their differences are defined as $\Delta\lambda$ and $\Delta K$.
+When the critical point $T_{p}=T_{c}$, we have $\Delta\lambda=\Delta K=0$.
+We will fit the numerical coefficients of $\Delta\lambda$ and $\Delta K$ according to the following form Wei and Liu (2018)
+$\frac{\Delta\lambda}{\lambda_{c}},\frac{\Delta K}{K_{c}}\sim\alpha(t-1)^{\beta}.$ (30)
+Here, $t=T_{p}/T_{c}$, $\alpha$ and $\beta$ are real numbers, the subscript “cc” denotes the value of the Lyapunov exponent and Gaussian curvature at the critical point.
+Figure 3: The reduced null Lyapunov exponents and Gaussian curvatures for the coexistence small and large black holes.
+(a) Null Lyapunov exponents and (b) Gaussian curvature $K_{HL}$ versus reduced temperature $t$ for a Hayward-Letelier-AdS black hole.
+As shown in Fig.
+3, we plot the normalized differences ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta K}/{K_{c}}$ as functions of the reduced temperature $t$.
+When $T_{p}=T_{c}$, ${\Delta\lambda}/{\lambda_{c}}={\Delta K}/{K_{c}}=0$.
+When $T_{p}$ deviates from $T_{c}$, both ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta K}/{K_{c}}$ take non-zero values, indicating the emergence of discontinuity.
+Ultimately, we obtain the following analytical expressions
+$\frac{\Delta\lambda}{\lambda_{c}}\simeq 6.9374(t-1)^{0.5761},$ (31)
+and
+$\frac{\Delta K}{K_{c}}\simeq 49.5398(t-1)^{0.6032}.$ (32)
+These results indicate that ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta K}/{K_{c}}$ can serve as order parameters.
+However, their critical exponents deviate from the exponent of $1/2$ observed in Reissner-Nordström black holes, suggesting that the geometry and dynamics of regular black holes exhibit richer critical behavior compared to singular black holes.
 V Discussion
 In this work, we systematically investigate the profound connections between spacetime geometry, thermodynamics, and chaos dynamics during first-order phase transitions of black holes.
-We begin by reviewing the behavior of first-order phase transitions of black holes as thermodynamic systems, noting the limitations of conventional approaches that rely on thermodynamic potentials.
-We then introduce the Gaussian curvature $K$ as an intrinsic geometric quantity, and using a known relation $K(r_{LR})=-\lambda^{2}(r_{LR})$ for unstable null orbits, we link intrinsic geometry to dynamical chaos and black hole phase transitions.
-Through numerical analysis of three spherically symmetric black holes, the RN-AdS, the Hayward-AdS, and the Hayward-Letelier-AdS models, we find that when system parameters $\tilde{Q}$ and $\tilde{g}$ enter the spinodal region, the free energy exhibits a swallowtail structure, while both the Gaussian curvature $K$ and the Lyapunov exponent $\lambda$ display multivalued behavior.
+While the standard thermodynamic methods describe phase transitions through the behavior of thermodynamic potentials, the corresponding evolution of spacetime geometry itself demands a more direct description.
+Based on this consideration, we examine the possibility of employing the intrinsic geometric quantity Gaussian curvature $K$ as a new probe.
+We find that during a first-order phase transition, the Gaussian curvature of unstable null orbits exhibits multivalued behavior inside the spinodal region and coincides precisely with the swallowtail structure of the free energy.
+This phenomenon can be understood through the known relation $K(r_{LR})=-\lambda^{2}(r_{LR})$, whereby the Gaussian curvature inherits the multivalued nature of the Lyapunov exponent.
+However, this relation should not be interpreted as a mere translation of the dynamical results into geometric language.
+The key advance of our work lies in establishing a self-contained geometric diagnostic framework that operates independently of any dynamical calculation.
+The procedure is purely geometric: first, determining the light ring radius from the geodesic curvature of the optical metric, then computing the corresponding Gaussian curvature via the first fundamental form, and finally, checking for multivalued behavior.
+This self-consistent procedure yields results in precise agreement with thermodynamic predictions.
+Through numerical analysis of the Hayward-Letelier-AdS black hole, we find that when the parameter $\tilde{g}$ enters the spinodal region, the free energy exhibits a swallowtail structure, while both the Gaussian curvature $K$ and the Lyapunov exponent $\lambda$ exhibit multivalued behavior.
 The temperature intervals over which this occurs align precisely with the thermodynamic spinodal region.
-Near the critical point, both $K$ and $\lambda$ diverge with a critical exponent of $1/2$, following the scaling relation $(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}$, consistent with mean-field theory.
 Furthermore, we demonstrate that in the absence of a phase transition ($\tilde{g}>\tilde{g}_{c}$), the Gaussian curvature $K$ varies monotonically and the geometric multivaluedness disappears.
-These results provide the first evidence of spacetime degeneracy accompanying black hole phase transitions and indicate that Gaussian curvature can serve as a diagnostic tool for phase transitions independent of thermodynamic potentials.
-This opens a new pathway for studying black hole phase structures from a purely geometric perspective.
-Appendix A: Potential approach
-For a spherically symmetric metric
-$ds^{2}=-f(r)dt^{2}+\frac{1}{g(r)}dr^{2}+r^{2}d\Omega^{2},$ (47)
-the Lagrangian reads
-$\mathcal{L}=g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=\sigma.$ (48)
-When $\theta=\frac{\pi}{2}$, we obtain the equation of motion
-$(\dot{r})^{2}=V_{\text{eff}}(r).$ (49)
-The effective potential is given by
-$V_{\text{eff}}(r)=g(r)\left[\frac{E^{2}}{f(r)}-\frac{L^{2}}{r^{2}}-\sigma\right],$ (50)
-where $E$ and $L$ are defined by Eq.(14). For massless particles in circular orbits, $\sigma=0$, and the unstable circular orbit is denoted as $r_{LR}$. We can derive from
-$V_{\text{eff}}(r_{LR})=0,\quad V^{\prime}_{\text{eff}}(r_{LR})=0.$ (51)
-This leads to
-$f^{\prime}(r_{LR})=\frac{2f(r_{LR})}{r_{LR}}.$ (52)
-For massive particles in circular orbits, $\sigma=1$, and the unstable circular orbit is denoted as $r_{0}$. The orbit satisfies
-$V_{\text{eff}}(r_{0})=0,\quad V^{\prime}_{\text{eff}}(r_{0})=0.$ (53)
-This results in
-$L^{2}=\left(\frac{f^{\prime}r^{3}}{2f-rf^{\prime}}\right)_{r=r_{0}},\quad E^{2}=\left(\frac{2f^{2}}{2f-rf^{\prime}}\right)_{r=r_{0}}.$ (54)
-Appendix B: Geometric approach
-The light ring $r_{LR}$ can also be determined independently using a geometric approach.
-Consider the optical metric given in Eq. (7). Its geodesic curvature $\kappa_{g}$ is given by
-$\kappa_{g}=\sqrt{\frac{g(r)}{f(r)}}\frac{2f(r)-rf^{\prime}(r)}{2r}.$ (55)
-On the light ring $r=r_{LR}$, the geodesic curvature vanishes, yielding
-$2f(r_{LR})=r_{LR}f^{\prime}(r_{LR}).$ (56)
-This result coincide with Eq. (52).
+This confirms that the multivalued behavior of the Gaussian curvature is a geometric signature of the phase transition.
+Near the critical point, the scaling laws of the Gaussian curvature and the null Lyapunov exponent for the Hayward-Letelier-AdS black hole deviate from those observed in most black hole solutions, suggesting that regular black holes may exhibit more interesting dynamical and geometric behaviors.
+These results advance the paradigm from standard thermodynamic or dynamical frameworks toward a purely geometric perspective.
+They indicate that black hole first-order phase transitions can be fundamentally understood as a branching behavior in the curvature structure of spacetime.
+Our work provides the direct evidence of the accompanying degeneracy in spacetime geometry during such phase transitions and demonstrates that Gaussian curvature can serve as a self-consistent diagnostic probe.
+This opens a new pathway for studying black hole phase structures from a geometric perspective.
 VI Acknowledgements